OpenStudy (anonymous):

what is x plus 4 over 3 equals 2

OpenStudy (anonymous):

x + \(\dfrac{4}{3}\) = 2

OpenStudy (anonymous):

Subtract 2 - \(\dfrac{4}{3}\)

OpenStudy (anonymous):

(x +4)/3 =2 x+4 = 6 x = 2